You are looking at a new project and have estimated the following cash flows, net income and book value data:

Year 0: CF = -165,000

Year 1: CF =    63,000

Year 2: CF =    70,000

Year 3: CF =    91,000

Your required return for assets of this risk is 12%. What is the NPV for this project?

Following the above problem, what is the payback period of this project? If the cutoff of your company’s payback period is 3 years, should you accept this project?
Answer #1

a.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=63000/1.12+70,000/1.12^2+91000/1.12^3

=176825.57

NPV=Present value of inflows-Present value of outflows

=176825.57-165,000

=11825.57(Approx).

b.

Year Cash flows Cumulative Cash flows
0 (165,000) (165,000)
1 63,000 (102,000)
2 70,000 (32000)
3 91,000 59000

=2+(32000/91000)

Payback period=Last period with a negative cumulative cash flow+(Absolute value of cumulative cash flows at that period/Cash flow after that period).

=2.35 years(Approx)

Hence since payback is less than 3 years;project must be accepted.